OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? 3x - y = -11 3x + y = - 7 3x + y = 5 x + 3y = 3 @iGreen.

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

well you would find the slope so 6/-2

OpenStudy (anonymous):

then divide so x=3

OpenStudy (anonymous):

No, 6 / -2 isn't 3..check again.

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Yep, so the slope is -3..

OpenStudy (anonymous):

Which options have slopes of -3? Convert them into the form of y = mx + b, where m is slope.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Like this: 3x - y = -11 Subtract 3x to both sides: -y = -3x - 11 Multiply -1 to all terms: y = 3x + 11 So what's the slope of this equation?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Yep, so we can eliminate this one, because it has a slope of 3.

OpenStudy (anonymous):

The answer is B

OpenStudy (anonymous):

y=-3x-7

OpenStudy (anonymous):

Anyway, to find the equation, we can plug in the slope with one of the points into point-slope form. \(y - y_1 = m(x - x_1)\) \(y - 2 = -3(x +3)\) Distribute -3 into the parenthesis: \(y - 2 = -3x - 9\) Add 2 to both sides: \(y = -3x - 7~\Large\color{lime} \checkmark\)